OpenStudy (anonymous):

Find the linear approximation to f(x)= (1)/(x^2 + 1) at x=1and use it to approximate f(1.1)

OpenStudy (anonymous):

The idea behind linear approximation is that you want to draw a line that is really close to the curve at a specific point so you can estimate the point on the curve. Since they told us to find the linear approximate of f(1.1), then the tangent line at f(1) will make a great starting point. remember y=mx + b is the equation for a line we have an (x) value, and we can plug it in to find the (y) value: f(1) = 1/2 so you're point will be (1, 0.5) Remember that the definition of a derivative is just the slope of the tangent line? well then all we have to do is find the derivative and we can write our linear equation out. \[f(x) = 1/(x^2 + 1) \] \[f'(x) = -2x/(x^2 + 1)^2\] lets plug in the x=1 to find the slope of our line \[f'(1) = -2/4 = -1/2 = -0.5\] so our equation is: y= -0.5 (x) + b to find b, we can plug in our ( 1, 0.5) 0.5 = -0.5 + b b = 1 so our linear equation is: y = -0.5 (x) + 1 now we can just plug in our x=1.1 to find f(1.1) y = -0.5 (1.1) + 1 y = -0.55 + 1 y = 0.45 so the linear approximate of f(1.1) is about 0.45